LEYDEN – AMSTERDAM – COLUMBIA

SUMMER PROGRAM IN AMERICAN LAW

June 30 – July 26, 2002

The Law Faculties of Columbia University in the City of New York, Leiden University, and the University of Amsterdam announce that the fortieth Summer Program in American Law will be offered at Leiden University from June 30 through July 26, 2002.

Since 1963 these summer courses are held alternately at the Leiden Law School and at the Amsterdam Law School and are designed to provide a general introduction to the American legal system with emphasis on areas of particular interest to European lawyers.

Preface

Go west, young man! Once upon a time (not in the west, of course, but in the east), this slogan was

introduced to stimulate young Americans to pick up their things and move towards the western frontier. Nowadays, this slogan is more likely to raise all sorts of interesting questions, particularly between smart lawyers. Think, for instance, about the a contrario reasoning: may a young man also go east? And what about the equality of the sexes: is a young woman also welcome in the west, and respectively the east? But let me not get too lawyerish in this short preface. After all, I only use the slogan because it summarizes, at least partially, what the Leyden- Amsterdam- Columbia Summer Program is all about. You, the reader of this brochure, whether young or old, man or woman, are being offered a unique opportunity to study a legal system from - for most of us - the west, i.e., the American legal system. On the other hand - and very conveniently so, while your mind is traveling westwards, your body is permitted to stay in the east.

The Leyden-Amsterdam-Columbia Summer Program in American Law started in 1963, has been organized ever since and is therefore, in 2002, in its 40th year. For more than a generation now, each month of July a group of professors from the Columbia University School of Law, situated in New York City, travels to Leiden or Amsterdam to teach courses on American law. Both the contents of these courses and the teaching method used is the “all-American way”. As such, both contents and teaching method will be new and fascinating for you. In four weeks you will become acquainted with a selection of interesting and important fields of American law. Moreover, you will (be forced to) get used to the so-called Socratic method: each day the instructor will assign a bunch of cases and materials to be studied for next day, and next day the instructor will call upon you to or your neighbor to discuss the assignments in class. So be prepared! And do not complain that I did not warn you!

After a day or a week of hard work there must be an opportunity to relax. The Summer Program takes care of that too. Various excursions in and outside Leiden will be organized, during weekdays and in the weekends. And when nothing is on the program and the assignments have been studied, you will form a small group with a few other participants to discover what Leiden has to offer by way of terraces, discos and beaches. Living there myself, I may add that there is quite a lot to discover.

This preface is not meant as yet another piece of advertisement, either for the Summer Program or for the City of Leiden. The rest of this brochure should be sufficiently tantalizing to convince you that participation is the thing to do and that your money will be invested well. Speaking of money, the Summer Program would not have achieved the age of 40 without the sponsors mentioned in this brochure. The Board of Directors is very grateful for this financial support. But, as we all know, there is never enough money. Therefore, the Board hopes that other sponsors will join the Summer Program, in particular the alumni of the Summer Program. Alumni are, after all, the perfect witnesses of the Summer Program’s success. In this you can believe me, because I am an alumnus myself.

One last word: should you, the reader of this brochure, not be interested in our Summer Program - perhaps you already booked a holiday in July or you are also a (sponsoring) alumnus? - you know better than to throw it away! Give it to someone who might be interested. And start spreading the news: registration for the Leiden session 2002 has started.

Maurice V. Polak Board of Directors

Introduction

Since 1963 the faculty of Columbia Law School, in cooperation with our partners in the law faculties of

Leiden and Amsterdam, have offered this Summer Program in American Law to introduce our legal system to students from the world over. We have learned a great deal in forty years about how to present a combination of basic material and advanced legal analysis to young lawyers from other systems. They, in turn, have taught us a great deal about our craft of teaching and, through their questions and discussions, about our law.

This coming summer my colleagues will return to Leiden with an extraordinary program of offerings. In addition to our required introductions to constitutional law, civil litigation, statutes and administrative practice, we will present a newly redesigned introduction to researching American law. This course, which has been under development for several years and has been refined through presentation in a variety of settings, offers what we at Columbia believe is the most effective route to rapid mastery of the materials of our law. Through explanatory lectures and hands-on training we can afford students and practitioners who need the ability to find US law a range of skills that we think are available nowhere else in a short course. For those considering an LLM degree in a US law school, whether at Columbia or elsewhere, we believe these offerings provide an invaluable foundation to make your LLM study more successful and rewarding.

Along with this expanded basic curriculum, we will present a wide range of elective courses exploring fastdeveloping areas of US law. From the fundamentals of contracts law to to the legal revolution sparked by the Internet, scholars studying and participating in current legal change will be bringing you the newest developments, and the context in which to understand them. Students and young practitioners interested in these areas will find both a comprehensive survey of US law as it stands, and also an opportunity to discuss forthcoming developments with professors who are among those best placed to observe and affect them.

One of the unique features of American legal education has always been the give-and-take of classroom discussion. My colleagues from the Columbia faculty volunteer to participate in the Program because of the excitement it offers them as teachers, giving them a chance to bring their style of class discussion to a different and fascinating range of students. We urge you to come to meet us in Leiden in July 2002, to be part of a Program that is in its second generation of achieving extraordinary results, for participants and faculty alike.

Eben Moglen Executive Director

Program of Instruction

To assure a shared basic knowledge, Statutes, Civil Procedure and Constitutional Law are compulsory courses for all participants. In addition, each participant is required to enroll in at least three elective courses in Legal Research, Contracts, Human Rights, Internet Law, or Criminal Law and Process.

As the number of participants in an elective course may be limited, applicants are requested to list on the registration form all available electives in order of their preference.

Compulsory Courses

Civil Procedure Dean Ellen Chapnick

This course provides an overview introduction of the private-law litigation system in the United States. Primary topics are: the structure and interrelation of U.S. state and federal courts, procedural mechanisms of litigation control, and the unique civil procedure associated with the distinctively American institution of the civil jury trial. Intensive attention will be paid to the analytical process of reading American case decisions.

Statutes & Regulations, Environmental Law Professor Edward Lloyd

This course will consider some fundamental structural characteristics of the American political and legal system having particular importance for public law, with specific application to the field of environmental law. We will examine American approaches to the materials of public policy: statutes, regulations, and the institutions that create and administer them. Topics will include Congress, the President and the administrative agencies: the process for forming statutes and regulations and current disputes about the proper materials and techniques of interpretation. The context of American environmental regulation illuminates the complex interplay of institutions and interpretations that results in the uncertain implementation of social policy.

Constitutional Law Staff

This is the basic course in constitutional law, a foundation for more specialized courses on the Constitution and for public law courses generally. The course locates the Constitution in the life of the United States. It explores: the theory of the Constitution and its antecedents, judicial review, its justification and development, and its legal and political significance; the nature of our federal system, the growth of national power and of limitations on state authority, and the abiding significance of the states; the separation of the powers and varieties of checks and balances in the U.S. government; and the theory and content of individual rights under the Constitution, the development of the principal rights during 200 years by Constitutional amendment and judicial interpretation, and the jurisprudence of the Judiciary in its role as the guardian of rights under the Constitutions and civil rights act.

Elective Courses

Human Rights in International Constitutional Perspective Professor Jack Greenberg

The course will compare, and build bridges between, the law of International Human Rights and the Constitutional law of rights in the United States. We will examine the history, theory, and sources of rights in the two systems, the institutions for implementing them, and the remedies for violations. We will also compare the treatment of particular rights in the two systems: for example, the right to life (capital punishment?); freedom of expression (including hate speech); freedom of religion; equality and the equal protection of the laws (including affirmative action); and economic and social rights. Topics will be considered in comparison to laws of other countries.

Contracts Staff

This course is an introduction to the central themes of American contract law and a presentation of modern issues affecting contracts. Special attention is paid to those aspects of contract law that distinguish U.S. contracts law from the commercial law systems of West Europe.

Internet Law Professor Eben Moglen

This course considers the legal significance of the development of computer-assisted communications, including the network of computer networks known at the Internet. American law has begun to adjust to conditions created by technological changes that the rest of the world will experience in the next decade. The goal here is to present the American experience, along with some social theory helpful in understanding the relation between rapid technological change and the legal system’s response. Topics covered will include the law of encryption, secrecy and anonymity; the effect of computer-assisted communication on contracts rules and private international law; new challenges to the visibility of the intellectual property system; and the application of competition law to the new technological environment. No prior experience with the Internet is required, but students will get a quick hands-on introduction to some of the technology in question.

Legal Research — Beginning to Advanced Instruction in Finding U.S. Law Kent McKeever

This course uses print and electronic research practices to introduce participants to the textual sources of American law. Focusing on the processes that produce statutes, case reports and regulations, as well as their dissemination in print and electronic forms, the course shows how to become a proficient researcher in American legal material. Students will learn the basic tools of the practicing attorney and the skills of finding and merging texts in particular situations to provide an accurate basis for analysis and decision-making. The course will also review lawmaking and publishing from a comparative viewpoint, so that the distinctive requirements of American legal research can be emphasized.

Federal Criminal Law and Enforcement, and Its International Dimensions Professor Daniel Richman

This course will provide an overview of the American criminal justice system, with particular attention to the Federal Government’s enforcement bureaucracy and authority. Among the topics to be explored are the formal adjudicatory process and the negotiated dispositions that usually substitute for it; some basic federal criminal statutes; relations between federal and local enforcers, and between federal enforcers and international authorities; and responses to international terrorism.

Schedule of Classes

Week 1^{a}  Week 2  Week 3^{b}  Week 4^{c}
July 1  July 8  July 15  July 22
09.30 - 10.30 Civ. Pro.  Civ. Pro.  Crim. Law  Crim. Law
11.00 - 12.00  Statutes  Statutes  Human Rights Human Rights

LUNCH

2.00 - 3.00  Research  Research  Contracts  Contracts
3.30 - 4.30  Internet Law Internet Law Const’l Law  Const’l Law

aOn Monday, July 1, 2002, there will be a reception, starting at 7:00 p.m.

bOn Wednesday July 17, 2002, a simulated jury trial will be conducted.

cOn Friday July 26, 2002, a farewell dinner will conclude the Program.

Faculty

Ellen Chapnick Assistant Dean for Public Interest Programs

Assistant Dean Ellen P. Chapnick is the founding director of the Center for Public Interest Law at Columbia Law School. She joined Columbia after 20 years as a federal litigator including as the senior partner responsible for the environmental law department at Wolf Pepper Ross Wolf & Jones where, among other matters, she served as a plaintiff’s lawyer in the Exxon Valdez Oil Spill Litigation, for which she and her co-counsel shared the Trial Lawyers for Public Justices 1995 Trial Lawyer of the Year Award. Dean Chapnick also was a staff attorney for federal

Her recent pro bono work includes being co-president of the Center for Constitutional Rights, co-chairing the Court as Employer Subcommittee of the Second Circuit Task Force on Gender, Racial and Ethnic Fairness, and serving on the Association of American Law Schools Task Force on Pro Bono and Public Service Opportunities, of which she is the chair-elect. Dean Chapnick has received the 1997 Pro Bono Publico Award from Pro Bono Students America and a 1998 award in recognition of outstanding pro bono publico service from the Legal Aid Society of the City of New York. She is the author of several articles and the Access to the Courts chapter in the ABA’s The Law of Environmental Justice.

Dean Chapnick is an honors graduate of the Georgetown University Law Center and Cornell University, College of Arts and Sciences.litigation at the Puerto Rican Institute for Civil Rights in Puerto Rico and various labor unions.

Jack Greenberg Professor of Law

B.A., Columbia, 1945; LL.B., 1948; LL.D., Central State College (Ohio), 1965; Morgan State College, 1965; Lincoln (Pennsylvania), 1977; John Jay College of Criminal Justice, 1983; Columbia, 1984; De Paul, 1994. Assistant counsel, NAACP Legal Defense and Educational Fund, 1949-61; director-counsel, 1961-84. Argued before U.S. Supreme Court in forty cases, including Brown v. Board of Education, 1954, which declared “separate but equal” unconstitutional; argued other cases in the Supreme Court and other courts involving civil rights in all its aspects. Founder, Earl Warren Legal Training Program. Participated in human rights missions to the Soviet Union, Poland, South Africa, the Philippines, Korea, Nepal, and elsewhere. Adjunct professor, Columbia Law School, 1970-84; visiting lecturer, Yale Law School, 1971; Harvard Law School, 1983. Dean, Columbia College, 1989-93. Visiting professor, College of the City of New York, 1977; University of Tokyo Faculty of Law, 1993-94; distinguished visiting professor, St. Louis University Law School, 1994. Visiting professor, Lewis and Clark Law School, 1994, 1996; Princeton University, 1995; and University of Munich, 1998; Tokyo University, 1996, 1998; University of Nuremberg-Erlangen, 1999-2000. Cardozo lecturer, Association of the Bar of the City of New York, 1973. Fellow, American Academy of Arts and Sciences, 1998. Awarded American Bar Association Thurgood Marshall Award, 1996. Founding member, Mexican-American Legal Defense and Education Fund. Member, boards of the Asian American Legal Defense and Education Fund; Human Rights Watch (1978-98); and NAACP Legal Defense and Educational Fund. Awarded the Presidential Citizens Medal (2001). Publications include Race Relations and American Law (1959); “Litigation for Social Change,” (1973); Cases and Materials on Judicial Process and Social Change (1976); Dean Cuisine: The Liberated Man’s Guide to Fine Cooking (with Vorenberg, 1991); Crusaders in the Courts: How a Dedicated Band of Lawyers Fought for the Civil Rights Revolution (1994); and articles on civil rights, capital punishment, and other subjects. Teaching interests include constitutional, civil rights, and human rights law, civil procedure, and South Africa’s post-apartheid constitution.

Edward Lloyd Evan M. Frankel Clinical Professor in Environmental Law

B.A., chemistry, Princeton, 1970; J.D., University of Wisconsin, 1973; served as staff attorney and executive director of the New Jersey Public Interest Research Group, 1974-83. Currently serving, since 1983, as general counsel to the New Jersey Public Interest Research Group. Founding director of the Rutgers University Law School Environmental Law Clinic, 1985-2000. Lectured on environmental legal issues at Judicial College for New Jersey judges, on citizen suit litigation at the Practicing Law Institute, and on numerous environmental courses for the practicing bar at the New Jersey Institute for Legal Education. Currently co-founder and codirector of the Eastern Environmental Law Center and a member of the Litigation Review Committee of the Environmental Defense Fund, 1991-present. Served as member, New Jersey Supreme Court Committee on Environmental Litigation; conferee, Governor’s Conference on Electricity Policy, Planning and Regulation; and chairman, board of directors, New Jersey Environmental Voters’ Alliance, 1983-87. Testified before the U.S. Senate and House of Representatives committees on environmental enforcement.

Kent McKeever Director Columbia Law School Library

Kent McKeever has been the Director of the Arthur W. Diamond Law Library of the Columbia Law School since 1996. Before that he had served as the Acting Law Librarian, the Head of Technical Services, and the International, Foreign, and Comparative Law Librarian at Columbia, and as a Reference Law Librarian at the Fordham Law School Library. He received his Juris Doctor and his Masters in Library Science simultaneously from Lousiana State University in 1980. He has also served as a Lecturer at the Columbia University School of Library Service from 1986 through 1991, and has taught in a variety of professional development programs ranging from Fudan University in Shanghai to a US AID program in Almaty, Kazakhstan.

Eben Moglen Professor of Law Executive Director

B.A., Swarthmore, 1980; J.D., M.Phil., Yale, 1985; Ph.D., 1993. Articles editor, Yale Law Journal. Law clerk to Judge Edward Weinfeld, Southern District of New York, 1985-86, and to United States Supreme Court Justice Thurgood Marshall, 1986-87. Joined the Columbia faculty, 1987. Principal areas of interest are Anglo-American legal history, constitutional law, computers and free expression, and intellectual property. Publications include Anarchism Triumphant: Free Software and the Death of Copyright, First Monday, August 1999; The Invisible Barbecue , 97 Colum.L.Rev. 945 (1997); Considering Zenger: Partisan Politics and the Legal Profession in Colonial New York , 94 Colum.L.Rev. 1495 (1994).

Daniel Richman Visiting Professor, Columbia University School of Law (Fall 2002) Professor, Fordham University School of Law (1992- )

B.A., Harvard College, 1980; J.D., Yale, 1984. Notes editor, Yale Law Journal. Law clerk to Chief Judge Wilfred Feinburg, Second Circuit Court of Appeals, 1984-85, and to United States Supreme Court Justice Thurgood Marshall, 1985-86. Assistant United States Attorney, U.S. Attorney’s Office for the Southern District of New York, 1987-92, and Chief Appellate Attorney, 1991-92. Visiting Associate Professor University of Virginia School of Law. 1996-97. Consultant to U.S. Department of Justice, Office of the Inspector General, 1997-2000; U.S. Treasury Department, Waco Administrative Review, 1993-1994. Principal areas of interest are Federal Criminal Law, Criminal Procedure, and Evidence.

General Information

Admission Requirements

Eligible for the Summer Program are law graduates, who are professionally active in the practice of law, industry, commerce, government, international organizations or related activities. Advanced undergraduate law students may be admitted in exceptional cases.

Applications for admission and scholarships are invited before May 30th, 2002:

Leiden University, Faculty of Law Columbia Summer Program Attn: Mrs. B. Zaaijer P.O. Box 9500 2300 RA Leiden, The Netherlands

fax: +31.71.5277298

e-mail: columbiasc@law.leidenuniv.nl

Language and Method of Instruction

Proficiency in English is required of all participants. Classes will be conducted in English, in the manner customary at the Columbia Law School. Participants will be called upon in class to discuss the materials previously assigned to them. In describing in your application your schooling and proficiency in the English language, you are requested to be as specific as possible and to give a frank evaluation of your ability to study and discuss American legal materials. Your objectivity in this respect will avoid subsequent disappointment. Having passed a TOEFL test is one of the ways in which you can indicate your proficiency in English.

Classes

Participants will receive due notice of any changes in the curriculum that may prove to be necessary. Each participant is required to take courses in Constitutional Law, Statutes & Regulations and Civil Procedure, and at least three elective courses. Since the number of participants in each elective course may be limited, applicants are requested to state four elective courses in order of their preference. Ten hours of classroom discussion are scheduled for each course. Participants are required to attend all classes in the courses for which they have registered. Adequate preparation is expected of each participant. Upon completion of the Program, a certificate of attendance will be issued to all participants who have met these requirements. There will be no final examination.

Attendance

Attendance of classes is compulsory. Failure to attend classes will result in withdrawal of the right to receive the certificate, at the discretion of the Board of Directors.

Study materials

In most courses, participants will be provided with casebooks of the kind used in American law schools. In some courses, only mimeographed materials will be distributed. Study materials will be distributed upon registration on Sunday, June 30, 2002.

Preliminary reading

Participants are advised to read: A. Farnsworth, An Introduction to the legal System of the United States (Oceana Publications, New York) before attending the Program..

Reading room and libraries

At the faculty building, where all classes will be held, a special reading room is available to all participants. Participants may consult the American law collection.

Expenses

Fee

A fee of (Euro) E 1500 covers tuition, study materials for six courses, and all administrative expenses, including those of the extra curricular activities and the farewell dinner. Participants who have not yet graduated will pay a reduced fee of (Euro) E 750. Meals and lodgings will be charged separately.

Lodgings

All participants reside in student houses in Leiden. Under no circumstances will payment for lodging be refunded. A price of about (Euro) E 200 is charged for the duration of the Program.

Meals

On days on which classes are given, all participants are expected to attend the lunches organized by the Program. The total amount charged for 20 lunches is approximately (Euro) E 225. As a rule, friends or relatives cannot be admitted to the lunches. No other meals will be provided. Upon registration each participant will be supplied with a list of suitable restaurants in Leiden.

The total amount charged for tuition, lunches and lodging will be approximately E 1925 (students E 1175).

Scholarships

A limited number of scholarships is available for participants unable to raise the necessary funds on their own or with the help of their employer, family or friends. The scholarships may cover a) tuition, b) lodging,

c) lunches, or d) combinations of these, but on no account do they cover travel or personal expenses. When applying for a scholarship, please indicate the absolute minimum amount necessary to enable you to participate, and the reasons why you have no alternative but to apply to the Program for financial assistance. Applicants eligible for scholarships under arrangements between their respective governments and the government of the Netherlands must address themselves directly to the appropriate authorities in their own country.

Moreover, please note that students from certain countries (e.g. Germany) may apply for scholarships in accordance with arrangements between their respective governments and the government of The Netherlands. In such cases applicants must address themselves directly to the appropriate authorities of their own country.

Hans Smit Scholarship

In honor of Hans Smit, Stanley H. Fuld Professor of Law, Columbia University, who initiated the Leyden-Amsterdam-Columbia Summer Program in American Law in 1963, a special scholarship has been established. Each year one promising candidate will be selected who would not be able to attend without financial assistance. The Hans Smit Scholarship covers tuition, lodging, meals, and pocket money.

Administrative matters

Arrival and registration

It is imperative that all participants arrive on Sunday, June 30, for the purpose of registration and allocation of housing, between 12.00 - 6.00 p.m. The location of the registration desk will be announced in due course. In the evening from 8.00 p.m. onwards an informal gathering will take place for the participants, the Directors of the Summer Program, faculty members and staff. All participants are requested to attend this meeting.

Inaugural session

The inaugural session will take place on Monday, July 1, 2002, from 7:00 p.m. on. The inaugural session will be followed by a reception for participants, faculty members, and invited guests.

Extra Curricular activities

(Subject to change)

Saturday, July 6 Excursion Friday, July 12 Excursion Wednesday, July 17 Moot Court Saturday, July 20 Excursion Friday, July 26 Farewell dinner

Directors of the Leiden Session

During the Program Professor Simons may be consulted by special appointment to be made through his secretary, tel. 071-527-7820. Professor Moglen may be consulted on all matters concerning the instructional part of the Program before June 30, 2002 via e-mail: moglen@columbia.edu or during the Program after his lectures.

Office of the Program

During the course, an information desk will be available. Here, participants can consult Mrs. Brigid Zaaijer, coordinator, and Ms. Sanne Bruinstroop, assistant, on all matters pertaining to the course.

Before July 1, Mrs. Zaaijer and Ms. Brunistroop may be consulted at +31.71.5277632; fax +31.5277298 or e-mail columbiasc@law.leidenuniv.nl

For information regarding the 2002 Leiden session, please refer to:

Leiden University, Faculty of Law

Columbia Summer Program Attn. Mrs. B. Zaaijer, LL P.O. Box 9500 2300 RA Leiden, The Netherlands

tel. +31.71.527.76.32 fax: +31.71.527.7298 email: columbiasc@law.leidenuniv.nl

Columbia Summer Program Leiden 2002

Executive Directors

W. Simons, Professor of Law Leiden Universtiy E. Moglen, Professor of Law Columbia University

Coordinator

Mrs. B. Zaaijer

Board of Directors

G.A. Bermann (Columbia) Th.M. de Boer (Amsterdam) E. Moglen (Columbia) M.V. Polak (Leiden) C. van Raad (Leiden) W. Simons (Leiden) K. Thomas (Columbia) Friedl Weiss (Amsterdam) J.W. Zwemmer (Amsterdam)